Problem Description: An ichthyologist wants to determine if a population of Lumpsuckers is evolving. He samples 5,000 fish and determines their genotype for an AFLP locus with two alleles: B (band) and Nb (no band). He notices that 3,500 individuals are homozygous for the B allele, with the rest of individuals equally divided amongst heterozygotes and Nb homozygotes.

Questions:

  1. How many individuals are there in the sample for the other two genotypes? (2 pts)

  2. What are the observed genotypic frequencies in the sample? (3 pts)

  3. What are the observed allelic frequencies in the sample? (4 pts)

  4. What are the expected genotypic frequencies under HWE for the sample? (6 pts)

  5. Is the population evolving? Explain (Hint: calculate the Χ2 value) (5 pts)
Answer #1

Answer: 1 mo of individuals with Heterogygous and Homo. zygous Recessive = 5000-3500 = 1500 . which are equally divided intoobserved frequency for recessive Allele = (mod Homogygous Recessive + 2 of Heterogy gous mo)/Total = (750+1/2 x 750)/5000 = (6 this population is evolving of Allelic frequency is shifting toward domimont type Means dominant Allele type is prefered of